two trains 120m & 80m in length are running in opposite direction with velocities 42kmph and 30kmph. In what time will they will completely cross each other?
Answer by mananth(16949)

Add the lengths of the trains = 200 = 0.2 km
add the speeds of the trains = 72
t= d/r
=0.2/72 hours
=0.00278 hours
=0.00278*3600
=10 seconds